OpenStudy (anonymous):

laplace inverse of 1/2(e^(-2x))((x^2-2x-1)/(x^3+x-2))

OpenStudy (nikvist):

Do you mean \[\mathcal{L}^{-1}\left\{\frac{e^{-2s}}{2}\frac{s^2-2s-1}{s^3+s-2}\right\}\]